OpenStudy (anonymous):

how would i solve this?Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5)

OpenStudy (anonymous):

y= -3x+20

OpenStudy (anonymous):

\[\frac{x-6}{6-5}=\frac{y-2}{2-5};\]\[x-6=-(y-2)/3\]\[y=-3x+20\]